# Entanglement engineering of optomechanical systems by reinforcement learning

### v1

@@ III.2 Superiority of deep RL-based control of entanglement engineering for benchmark control methods
Our deep RL, a model-free learning method, is implemented in the measurement-based feedback control framework for entanglement engineering in open quantum optomechanics. To appreciate its performance, we employ two benchmark methods for comparison: Bayesian [63, 43] and random control. To characterize the quantum-entanglement control performance, we use the following three quantities: $\langle E_{N}\rangle$, $\widetilde{E}_{N}$, and $\widetilde{R}$ in open quantum optomechanical systems with either linear or nonlinear interaction between the quantum cavity and oscillator modes. In particular, $\langle E_{N}\rangle$ is the logarithmic negativity averaged over ten successive episodes with a single environment, $\widetilde{E}_{N}$ is the corresponding average over one episode with $T=500$ time steps with a single quantum environment, and $\widetilde{R}$ denotes the ensemble-averaged value of the reward $R$ over a small number of multiple parallel quantum environments.
@@
Table 1: Results of entanglement engineering from deep RL-based, Bayesian, and random control. The observations are the expectation of the photon number $\langle\hat{n}_{p}\rangle$ and the WCM flow $\mathcal{I}(t)$ by homodyne detection, at the measurement rate $\eta=1$. The Bayesian hyperparameter is $\lambda_{opt}=10$ for the $\langle\hat{n}_{p}\rangle$ task and $\lambda_{opt}=2$ for the $\mathcal{I}(t)$ task. The action is the amplitude modulation of the laser, which is in the range $G\in[-5,5]\omega_{m}$. Displayed are the results of the average logarithmic negativity $\langle E_{N}\rangle/\log{2}$ with the standard deviation. For training and testing phases, $\langle E_{N}\rangle/\log{2}$ is averaged over ten end-training or testing episodes, each having $T=500$ time steps with the step size $dt=0.01\,\omega_{m}^{-1}$. Each observation is obtained by averaging over $\textnormal{ntraj}=1$ for $\langle\hat{n}_{p}\rangle$ and $\textnormal{ntraj}=5$ for $\mathcal{I}(t)$ with only one quantum environment through simulating the SME in Eq. (II.2), and ntraj denotes the number of independent trajectories from SME simulations.
Controller Condition $\langle\hat{n}_{p}\rangle$ ntraj$=1$ $\mathcal{I}(t)$ ntraj$=5$
@@
Figure 2: Performance in terms of $\langle E_{N}\rangle/\log 2$ over a long time interval, compared for deep RL-based, Bayesian, and random control methods with respect to two observable options: the expected value $\langle\hat{n}_{p}\rangle$ and the WCM photocurrent $\mathcal{I}(t)$. The deep RL controller is trained with $T=500$ time steps with the step size $dt=0.01\,\omega_{m}^{-1}$. For all three control methods, displayed are results from the testing phase for the following set of time steps: $T=[500,1000,1500,2000,2500,3000,3500,4000]$ at the measurement rate $\eta=1$. The decay rates of the cavity and of the mechanical mode are $\kappa=0.01\,\omega_{m}$ and $\gamma=10^{-2}\kappa$, respectively. The system is initialized in the pure state $|\psi\rangle=|10\rangle$. The figure conventions, which apply to this and all subsequent figures, are as follows. If the vertical axis is labeled as $\langle E_{N}\rangle/E_{0}$, it represents the normalized logarithmic negativity, with $E_{0}=\log 2$ as the target entanglement value. Otherwise, when the vertical axis is labeled as $\langle E_{N}\rangle$, $\widetilde{E}_{N}$, or $E_{N}(t)$, it represents the original value of the logarithmic negativity.
@@
Figure 3: Performance of deep-RL agent in the online training and testing phase. The characterizing quantities are the logarithmic negativity $E_{N}$ and the reward function $R$. The setting is: measurement rate $\eta=1$, initial state $|\psi\rangle=|10\rangle$, and the decay rate of cavity and mechanical modes $\kappa=0.01\,\omega_{m}$, $\gamma=10^{-2}\kappa$, respectively. (a,c) Performance measures in the online training phase, where the mean $\widetilde{E}_{N}$ is over one episode with $T=500$ time steps on the fifth quantum environment (only one environment) and the mean reward $\widetilde{R}$ is obtained from $\mathbb{N}=5$ parallel quantum environments. The solid traces are the moving-window average over $100$ episodes. (b,d) Performance measures during the testing phase, where the logarithmic negativity $E_{N}(t)$ and WCM flow $\mathcal{I}(t)$ are obtained with $T=4000$ time steps.
The benchmark control methods are briefly described, as follows. Bayesian control [63, 43] is a state-based feedback control of the stochastic process as governed by SME. In our case, the control law is given by
$G(t)=-\lambda|\langle\hat{n}_{p}\rangle(t)-0.5|\omega_{m}$
with $\langle\hat{n}_{p}\rangle(t)$ being the observation, where the hyperparameter $\lambda$ can be numerically optimized based on the performance. If the observation is $\mathcal{I}(t)$, the control flow will be in the form
$G(t)=-\lambda|\mathcal{I}(t)-0.5|\omega_{m},$
@@
in which the Wiener process blocks the performance to some degree. In Bayesian control, the smaller the variance in the measured current, the better the performance. For the random control method, the flow is generated by a uniform distribution in the action range $G\in[-5,5]\omega_{m}$. Note that the actions $G$ of random control and deep RL are in the same range while the one of Bayesian control is determined by the hyperparameter $\lambda$ and the state-based observation value or the WCM current. To make a fair comparison, $\lambda_{opt}$ is optimized within the action range $G\in[-5,5]\omega_{m}$. Specifically, the optimized hyperparameter $\lambda_{opt}$ corresponds to the best performance of Bayesian control in the set $\lambda\in\{1,2,\ldots,\lambda_{max}\}$, where $\lambda_{max}$ is the maximum integer of $\lambda$ to guarantee the action range $G\in[-5,5]\omega_{m}$.
@@
Table 1 displays the values of the averaged logarithmic negativity $\langle E_{N}\rangle/\log 2$ from the deep RL, Bayesian, and random control methods. For deep RL control, the PPO agent is trained with $T=500$ time steps but tested with a longer time horizon, e.g., $T=4000$ steps, as shown in Fig. 2. From the SME simulations [Eq. (II.2)], when the observation is the expectation of the photon number, the Bayesian control with the optimized hyperparameter outperforms the deep RL method. However, when the observation is the WCM flow, the deep RL control outperforms the Bayesian method. This is promising as the WCM current is directly experimentally accessible while the expected photon number is not. Regardless of the observation, random control is generally ineffective. Under deep RL control, it can also be noted that the performance with the observation of WCM current has a stable and smaller variance compared with that in the case where the observation is the expected photon number. The results by deep RL control from the observation of random measurement current tend to reduce the performance by about $20\%$ compared to the expected photon number. For Bayesian control, the reduction is about $40\%$. Overall, with the experimentally feasible observation of WCM, the deep RL controller stands out as the choice of entanglement control for quantum optomechanical systems.
III.3 Characterization of deep-RL-based control of entanglement engineering
@@
Figure 4: Effects of decay and measurement rates on the control performance. Shown are the values of the average logarithmic negativity for (a) decay rates $\kappa=[0,0.01,0.03,0.05]\omega_{m}$ with $\eta=0.5$ and (b) measurement rates $\eta=[0.05,0.1,0.3,0.5,0.7,1]$ with $\kappa=0.01\,\omega_{m}$. Because the quality of the optical cavity is generally lower than that of the mechanical oscillator [66], the rate setting is $\gamma=10^{-2}\kappa$. The error bars represent the standard deviation of the data points. The average operation is over ten end-training or testing episodes. The training and testing time steps are the same: $T=500$. The initial quantum pure state is $|\psi\rangle=|10\rangle$.
@@
We characterize the performance of our deep-RL-based control method in terms of the dissipation rate, measurement rate, and the randomness effect for the initial state. For the measurement rate $\eta=1$, the PPO agent is sequence-wise trained with the WCM flow. Figures 3(a) and 3(c) show the average logarithmic negativity $\widetilde{E}_{N}$ and the mean reward $\widetilde{R}$, respectively, versus the episode during the training phase, in which $\widetilde{E}_{N}$ and $\widetilde{R}$ are averaged over one and five parallel quantum environments, respectively. Both quantities ultimately converge due to the properly designed reward function $R(t)=-|\mathcal{I}(t)-0.5|$. Note that the variance of $\widetilde{E}_{N}$ is suppressed with the episodes, implying the mixture-robust nature of entanglement in the quantum optomechanical system. The testing phase is longer ($T=4000$ time steps) than the training phase ($T=500$ time steps) and the corresponding performance measures are shown in Figs. 3(b) and 3(d). In addition to the variance in the learning of the deep RL agent with the stochastic policy, the Gaussian Wiener process in the WCM current [Eq. (10)] and the stochastic collapse process stipulated by the SME [Eq. (II.2)] also contribute to the variances of the performance measures.
@@
Since the quantum optomechanical system is coupled to the vacuum bath, the coupling strength or disturbance between the classical and quantum environments will affect the control performance, as exemplified in Fig. 4(a). Previous experiments [66] demonstrated that the quality of the optical cavity is generally lower than that of the mechanical oscillator, i.e., $\gamma<\kappa$, so we set the decay rate of the oscillator at two orders of magnitude smaller than that of the cavity [65]: $\gamma=10^{-2}\kappa$. Figure 4(a) shows, for both the expectation and the measurement flow observations, the performances of the training and testing processes, which are consistent with each other in the sense that their mean values decrease and the variances increase with the decay rate. The origin of the performance fluctuations is the classical dissipation to the vacuum bath, which renders the system less controllable by laser.
@@
The uncertainty in the classical information extracted from the quantum system depends on the discrete-time step size $dt$ and the measurement rate $\eta$ [Eq. (10)], which directly determines the degree of the quantum-state stochastic collapse and quantum decoherence from the WCM term in Eq. (II.2). If the expectation of the photon number is the observation, the stronger the measurement rate (proportional to the measurement strength), the poorer the performance of deep-RL control as characterized by a decrease in the mean values and an increase in the uncertainties of $E_{N}$, as shown in Fig. 4(b), which originate from the intrinsic random process in the stochastic master equation induced by the measurement process. However, if the observation is the WCM flow, the weaker measurement rate will introduce larger variances in the observation signal and reduce the stochasticity of the process due to the incomplete/partial extracted information as described by the SME. In our case, the target mean value is on the order of $10^{-1}$, rendering necessary to introduce a Gaussian filter to reduce the uncertainty. The resulting performance of deep-RL control is approximately the same for $\eta\in[0.05,1]$, as shown in Fig. 4(b).
@@
Figure 5: Robustness of deep-RL method trained with pure or mixed states. (a) In the training and testing phase, performance of $\langle E_{N}\rangle/E_{0}$ for different initial mixed states (solid traces): $\rho=(1-p)|10\rangle\langle 10|+p|01\rangle\langle 01|$ with $p=[0,0.1,0.2,0.3,0.4,0.5]$. The dashed traces indicate the performances trained with random initial mixed states with the random variable $p\in[0,0.5]$. (b) Testing performance of two kinds of trained agents with $p\in[0,1]$: one trained with the pure initial state $|\psi\rangle=|10\rangle$ and another with random initial mixed states, which are distinguished by the color depth of the curve and the error bars. The blue and red curves denote the performances with the observation $\langle\hat{n}_{p}\rangle$ and $\mathcal{I}(t)$, respectively, with error bars. The decay rate of the cavity mode is $\kappa=0.01\,\omega_{m}$, the measurement rate is $\eta=0.5$, and the training and testing time steps are $T=500$.
Experimentally, mixed quantum states are more realizable than pure states due to the quantum decoherence with the classical environment, e.g., the vacuum bath. To address this issue, we assume that the initial state is a mixed state in the form of [90]
$\rho=(1-p)|10\rangle\langle 10|+p|01\rangle\langle 01|,$ (25)
where the parameter $p$ is fixed or a random variable $p\in[0,1]$. The beam-splitter Hamiltonian [Eq. (3)] stipulates that the photon and phonon modes are symmetric to each other, so $p$ can be rescaled into the interval $p\in[0,0.5]$. Figure 5(a) shows the performance with respect to the initial mixed quantum state with the same parameter $p$ for each training and testing episode (solid traces), where the complete mixed case with $p=0.5$ leads to the worst performance but still possesses entanglement to a significant extent. The reason lies in the inherent property of the beam-splitter Hamiltonian, which can create the maximum entangled states: $[|10\rangle+e^{i\varphi}|01\rangle]/\sqrt{2}$, with respect to the part of the quantum state, such as $|10\rangle$ or $|01\rangle$ through the linear interactions, regardless of whether it acts on a pure or a mixed state. In Fig. 5(a), the dashed traces display the performance during the training phase with a random initial mixed quantum state, which is generated by the random variable $p$ with the uniform distribution in the range of $p\in[0,0.5]$. The error bar characterizes the uncertainty over ten end-training episodes.
Figure 5(b) shows the testing performance of two kinds of trained models, one trained by the initial state $|\psi\rangle=|10\rangle$ and another by the random initial mixed state $\rho=(1-p)|10\rangle\langle 10|+p|01\rangle\langle 01|$ (distinguished by dark and light colors, respectively). Note that the beam-splitter Hamiltonian naturally creates a Bell state starting from either $|10\rangle$ or $|01\rangle$. Intuitively, the beam-splitter Hamiltonian will generate a mixed Bell states if the initial state is a classical mixture of $|10\rangle\langle 10|$ and $|01\rangle\langle 01|$. With the mixed probability $p=0.5$, it results in an equal mixture of the Bell states, as shown in Fig. 5. In the testing phase, the two trained models use the same initial state for a fixed value of $p$. The two models have a comparable performance, suggesting that the deep RL method is robust to the initial randomness in a mixed state. More specifically, during the testing phase, the observation is the expected photon number or WCM current. The worse performance occurs for $p=0.5$ and for other values of $p$, the performance is symmetric about $p=0.5$. The physical reason behind the symmetric performance lies in the symmetric role of the photon and photon modes in the beam splitter Hamiltonian. Note that the model trained with the observation being the measurement current displays a small difference in the performance measure [$\langle E_{N}\rangle$ over the whole probability interval $p\in[0,1]$] between the best and worst cases, with less uncertainties than the case where the observation is the expected photon number. Taken together, our deep-RL model trained by the weak measurement flow holds a lower mean performance but possesses robustness against mixed quantum states compared with the scenario based on observing the expected value of the photon number, due to the strong capability of RL in learning randomness and executing accurate high-dimensional data-fitting.
IV Entanglement engineering in nonlinear quantum optomechanics
Figure 6: Generating and stabilizing entanglement in a nonlinear open quantum optomechanical system by deep-RL measurement-based feedback control. (a) The RL agent acts in one or multiple parallel quantum optomechanical environments according to the parameterized policy and collects the following data: observation $O_{t}=\langle n_{p}\rangle(t)$ of the photon number, reward $R(t)=-|O_{t}-\langle n^{\textnormal{target}}_{p}\rangle(t)|$, and action $\Delta(t),\alpha_{L}(t)\in[-5,5]\omega_{m}$ of the driving laser for time $t$ flow in one episode with $T$ time steps. After each episode, the quantum optomechanical environment is reset. The policy is updated after one or several episodes by using minibatch data to maximize the accumulated reward, with the aim to achieve the desired entanglement between cavity-optical and mechanical modes as best as possible. (b) Photon and Phonon number distributions of the entangled states.

### v2

@@ Results
To characterize the quantum-entanglement control performance, we use the following three quantities: $\langle E_{N}\rangle$, $\widetilde{E}_{N}$, and $\widetilde{R}$ in open quantum optomechanical systems with either linear or nonlinear interaction between the quantum cavity and oscillator modes. In particular, $\langle E_{N}\rangle$ is the logarithmic negativity averaged over ten successive episodes with a single environment, $\widetilde{E}_{N}$ is the corresponding average over one episode with $T$ time steps in a single quantum environment, and $\widetilde{R}$ denotes the ensemble-averaged value of the reward $R$ over a small number of multiple parallel quantum environments for each episode. In our computations, all the control actions $G$, $\alpha_{L}$, or detuning $\Delta$, the nonlinear coupling $g_{0}$, and the dissipation coefficients $(\kappa,\gamma)$ are in units of $\omega_{m}$. The time unit is $\omega_{m}^{-1}$.
RL in linear quantum optomechanics
A quantum optomechanical system with linear photon-phonon interactions is governed by the beam-splitter Hamiltonian. In an optical experimental platform, a 50:50 beam splitter with the transformation angle $\pi/4$ can create an entangled Bell state between the two input optical modes [78, 79, 80]. Similarly, in a quantum optomechanical system, Bell states between photon and phonon modes can be realized by controlling the beam-splitter Hamiltonian. As a result, the maximally attainable value of the logarithmic negativity is $E_{N}\sim\log 2\sim 0.7$ (in the natural logarithmic base), corresponding to the maximally entangled Bell state, as shown in Fig. 1. This “best” entangled state can be realized by the model-free PPO agent, regardless of whether the observation is the expectation or WCM photocurrent. To see this, we note that, in the beam-splitter model, the initial quantum state is set as a pure state [81, 82]: $|\psi\rangle=|10\rangle$, where the photon is in the first excited mode and the phonon is in the vacuum mode. The partial observable of the quantum state for the PPO agent is set as the expectation of the photon number $\langle\hat{n}_{p}\rangle(t)=\langle\hat{P}_{1}\rangle(t)$ or the WCM photocurrent $\sqrt{\eta}\,\mathcal{I}(t)=\langle\sqrt{\eta}\hat{P}_{1}\rangle(t)+\frac{dW(t)}{\sqrt{4\eta}\;dt}$.
@@
Table 1: Results of entanglement engineering from deep RL-based, Bayesian, and random control. The observations are the expectation of the photon number $\langle\hat{n}_{p}\rangle$ and the WCM photocurent $\mathcal{I}(t)$ at the measurement rate $\eta=1$. The Bayesian hyperparameter is $\lambda_{opt}=10$ for the $\langle\hat{n}_{p}\rangle$ task and $\lambda_{opt}=2$ for the $\mathcal{I}(t)$ task. Displayed are the results of the average logarithmic negativity $\langle E_{N}\rangle/\log{2}$ with the standard deviation. For training and testing phases, $\langle E_{N}\rangle/\log{2}$ is averaged over ten end-training or testing episodes, each having $T=500$ time steps. Each observation is obtained by averaging over $n_{\textnormal{traj}}=1$ for $\langle\hat{n}_{p}\rangle$ and $n_{\textnormal{traj}}=5$ for $\mathcal{I}(t)$ through simulating the SME, and $n_{\textnormal{traj}}$ denotes the number of independent trajectories from SME simulations.
Controller Condition $\langle\hat{n}_{p}\rangle$ $n_{\textnormal{traj}}=1$ $\mathcal{I}(t)$ $n_{\textnormal{traj}}=5$
@@
Experimentally, directly measuring the entanglement, e.g., in terms of logarithmic negativity, for arbitrary entangled states is generally not viable. Identifying an experimentally feasible quantity to characterize the entanglement in arbitrary quantum systems remains challenging. We focus on the relationship between logarithmic negativity and the expected photon number, based on recent experiments on multiplexed photon number measurement [49, 83, 29, 68, 69, 84, 85]. To proceed, we note that the beam-splitter Hamiltonian is limited to a four-level basis, due to the following reasons: (1) only one energy level in the cavity mode of the initial state has been excited from the vacuum state, i.e., $|\psi\rangle=|10\rangle$, (2) the linear interaction serves only to transfer the quantum states between the cavity and mechanical mode (i.e., no quantum excitation), and (3) the system couples to the vacuum bath only at absolute zero temperature (i.e., without any thermal excitation), thereby blocking any interactions between higher-level quantum states. In this case, the maximum logarithmic negativity $E_{N}\sim\log 2\sim 0.7$ implies that the attained quantum state is the following Bell state
$|\Phi^{\varphi}\rangle=\frac{1}{\sqrt{2}}\left[|10\rangle+e^{i\varphi}|01\rangle\right],$
with the associated expected photon number $\langle\hat{n}^{\textnormal{target}}_{p}\rangle=\langle\hat{P}_{1}\rangle=0.5$. Consequently, the reward function can be set as $R_{t}\equiv-|O_{t}-0.5|$, regardless of whether the observation $O_{t}$ is $\langle\hat{n}_{p}\rangle(t)$ or $\mathcal{I}(t)$. Because of the relatively small target value of the expected photon number: $\langle\hat{P}_{1}\rangle=0.5$, the variance $\mathcal{I}(t)$ in the WCM photocurrent can be reduced by a Gaussian filter [86] with the weak measurement rate $\eta\leq 1$. The Gaussian kernel parameters of the filter such as the filter interval and the variance can be numerically chosen to reduce the standard deviation of the measurement photocurrent into a certain range, e.g., about ten times larger than the mean value (details in Supplementary Note 6). The PPO agent applies an updated stochastic policy to the quantum optomechanical environment to maximize the accumulated reward, where the action $G(t)$ is proportional to the amplitude of the cavity mode: $G(t)=g_{0}\bar{\alpha}_{c}(t)$. The action can be controlled by an incident laser [63] and is continuous in a certain range, e.g., $G\in[-5,5]\omega_{m}$. The decay rate of cavity and mechanical modes $\kappa=0.01\,\omega_{m}$, $\gamma=0.01\,\kappa$, respectively, because the quality of the mechanical oscillator mode is generally better than that of the optical cavity or microwave resonator mode [61, 59, 58, 65, 66, 67].
@@
Figure 2: Performance in terms of $\langle E_{N}\rangle/\log 2$ over a long time interval, compared for deep RL-based, Bayesian, and random control methods with respect to two observable options: the expected value $\langle\hat{n}_{p}\rangle$ and the WCM photocurrent $\mathcal{I}(t)$. The deep RL controller is trained with $T=500$ time steps. For all three control methods, displayed are results from the testing phase for the following set of time steps: $T=[500,1000,1500,2000,2500,3000,3500,4000]$ at the measurement rate $\eta=1$. The conventions, which apply to this and all subsequent figures, are as follows. If the vertical axis is labeled as $\langle E_{N}\rangle/E_{0}$, it represents the normalized logarithmic negativity, with $E_{0}=\log 2\sim 0.7$ (in the natural logarithmic base) as the target entanglement value. Otherwise, when the vertical axis is labeled as $\langle E_{N}\rangle$, $\widetilde{E}_{N}$, or $E_{N}(t)$, it represents the original value of the logarithmic negativity.
@@
Figure 3: Performance of deep-RL agent in the online training and testing phase. The characterizing quantities are the logarithmic negativity $E_{N}$ and the reward function $R$ with the measurement rate $\eta=1$. (a,c) Performance measures in the online training phase, where the mean $\widetilde{E}_{N}$ is over one episode with $T=500$ time steps on the fifth quantum environment (only one environment) and the mean reward $\widetilde{R}$ is obtained from $\mathbb{N}=5$ parallel quantum environments. (b,d) Performance measures during the testing phase, where the logarithmic negativity $E_{N}(t)$ and WCM photocurrent $\mathcal{I}(t)$ are obtained with $T=4000$ time steps. The solid traces represent the moving-window average over $100$ episodes for (a,c) and $100$ time steps for (b,d).
@@
Our deep RL, a model-free learning method, is implemented in the measurement-based feedback control framework for entanglement engineering in open quantum optomechanics. Details about the PPO algorithm applied in the linear quantum optomechanics are presented in Supplementary Note 4. To appreciate its performance, we employ two benchmark methods for comparison: Bayesian [54, 47] and random control. Bayesian control [54, 47] is a state-based feedback control of the stochastic process as governed by the SME. In our case, the control law is given by $G(t)=-\lambda|\langle\hat{n}_{p}\rangle(t)-0.5|\omega_{m}$ with $\langle\hat{n}_{p}\rangle(t)$ being the observation, where the hyperparameter $\lambda$ can be numerically optimized based on the performance. If the observation is $\mathcal{I}(t)$, the control flow will be in the form $G(t)=-\lambda|\mathcal{I}(t)-0.5|\omega_{m}$, in which the Wiener process blocks the performance to some degree. In Bayesian control, the smaller the variance in the measured photocurrent, the better the performance. For the random control method, the flow is generated by a uniform distribution in the action range $G\in[-5,5]\,\omega_{m}$. Note that the actions $G$ of random control and deep RL are in the same range while the one of Bayesian control is determined by the hyperparameter $\lambda$ and the state-based observation value or the WCM photocurrent. To make a fair comparison, $\lambda_{opt}$ is optimized within the action range $G\in[-5,5]\omega_{m}$. Specifically, the optimized hyperparameter $\lambda_{opt}$ corresponds to the best performance of Bayesian control in the set $\lambda\in\{1,2,\ldots,\lambda_{max}\}$, where $\lambda_{max}$ is the maximum integer of $\lambda$ to guarantee the action range $G\in[-5,5]\,\omega_{m}$.
@@
Table 1 displays the values of the averaged logarithmic negativity $\langle E_{N}\rangle/\log 2$ from the deep RL, Bayesian, and random control methods. From the SME simulations, when the observation is the expectation of the photon number, the Bayesian control with the optimized hyperparameter outperforms the deep RL method. However, when the observation is the WCM photocurrent, the deep RL control outperforms the Bayesian method. This is promising as the WCM photocurrent is directly experimentally accessible while the expected photon number is not. Regardless of the observation, random control is generally ineffective. The results by deep RL control from the observation of WCM photocurrent tend to reduce the performance by about $20\%$ compared to that based on the expected photon number. For Bayesian control, the reduction is about $40\%$. Moreover, Fig. 2 compares the long-time entanglement engineering for three control methods. Especially, for deep RL control, the PPO agent is trained with $T=500$ time steps but tested with a longer time horizon, e.g., $T=4000$ steps, including the unexplored regime by the PPO agent. It is worth noting from Fig. 2 that the performance of deep RL with the observation of WCM photocurrent exhibits a more stable and smaller variance compared to the case where the observation is the expected photon number, especially after $T=2000$. Overall, with the experimentally feasible observation of WCM, the deep RL controller stands out as the choice of entanglement control for quantum optomechanical systems.
@@
Figure 4: Effects of decay and measurement rates on the control performance. Shown are the values of the average logarithmic negativity for (a) decay rates $\kappa=[0,0.01,0.03,0.05]\,\omega_{m}$ with $\eta=0.5$ and $\gamma=0.01\,\kappa$, and (b) measurement rates $\eta=[0.05,0.1,0.3,0.5,0.7,1]$ with $\kappa=0.01\,\omega_{m}$ and $\gamma=0.01\,\kappa$. The error bars represent the standard deviation of the data points. The average operation is over ten end-training or testing episodes. The training and testing time steps are the same: $T=500$.
@@
We characterize the performance of our deep-RL-based control method in terms of the dissipation rate, measurement rate, and the randomness effect for the initial state. For the measurement rate $\eta=1$, the PPO agent is sequence-wise trained with the WCM photocurrent. Figures 3(a) and 3(c) show the average logarithmic negativity $\widetilde{E}_{N}$ and the mean reward $\widetilde{R}$, respectively, versus the episode during the training phase, in which $\widetilde{E}_{N}$ and $\widetilde{R}$ are averaged over one and five parallel quantum environments, respectively. Both quantities ultimately converge due to the properly designed reward function $R(t)=-|\mathcal{I}(t)-0.5|$. Note that the variance of $\widetilde{E}_{N}$ is suppressed with the episodes, implying the mixture-robust nature of entanglement in the quantum optomechanical system. The testing phase is longer ($T=4000$ time steps) than the training phase ($T=500$ time steps) and the corresponding performance measures are shown in Figs. 3(b) and 3(d). In addition to the variance in the learning of the deep RL agent with the stochastic policy, the Gaussian Wiener process in the WCM photocurrent and the stochastic collapse process stipulated by the SME also contribute to the variances of the performance measures. However, the deep RL control still manages to maintain the solid traces of the testing $\mathcal{I}(t)$ around the target value $\langle\hat{n}^{\textnormal{target}}_{p}\rangle=0.5$ in Fig. 3(d) and the resulting entanglement quantity $E_{N}(t)$ is displayed in Fig. 3(b).
@@
Since the quantum optomechanical system is coupled to the vacuum bath, the coupling strength or disturbance between the classical and quantum environments will affect the control performance, as exemplified in Fig. 4(a). Previous experiments [61, 64, 58, 59] demonstrated that the quality of the mechanical oscillator is generally better than that of the optical cavity or microwave resonator, i.e., $\gamma<\kappa$, so we set the decay rate of the oscillator at two orders of magnitude smaller than that of the cavity [56]: $\gamma=0.01\,\kappa$. Figure 4(a) shows, for both the expectation and the measurement flow observations, the performances of the training and testing processes, which are consistent with each other in the sense that their mean values decrease and the variances increase with the decay rate. The origin of the performance fluctuations is the classical dissipation to the vacuum bath, rendering the system less controllable by laser.
@@
The uncertainty in the classical information extracted from the quantum system depends on the discrete-time step size $dt$ and the measurement rate $\eta$, which directly determines the degree of the quantum-state stochastic collapse and quantum decoherence from the WCM term in the SME. If the expectation of the photon number is the observation, the stronger the measurement rate (proportional to the measurement strength), the poorer the performance of deep-RL control as characterized by a decrease in the mean values and an increase in the uncertainties of $E_{N}$, as shown in Fig. 4(b), which originate from the intrinsic random process in the SME induced by the measurement process. However, if the observation is the WCM photocurrent, the weaker measurement rate will introduce larger variances in the observation signal and reduce the stochasticity of the process due to the incomplete/partial extracted information as described by the SME. In our case, the target mean value, $\langle\hat{n}^{\textnormal{target}}_{p}\rangle=0.5$, is on the order of $10^{-1}$, rendering necessary introducing a Gaussian filter to reduce the uncertainty. The resulting performance of deep-RL control is approximately the same for $\eta\in[0.05,1]$, as shown in Fig. 4(b).
@@
Figure 5: Robustness of deep-RL method trained with pure or mixed states. (a) In the training and testing phase, performance of $\langle E_{N}\rangle/E_{0}$ for different initial mixed states (solid traces): $\rho=(1-p)|10\rangle\langle 10|+p|01\rangle\langle 01|$ with $p=[0,0.1,0.2,0.3,0.4,0.5]$. The dashed traces indicate the performances trained with random initial mixed states with the random variable $p\in[0,0.5]$. (b) Testing performance of two kinds of trained agents with $p\in[0,1]$: one trained with the pure initial state $|\psi\rangle=|10\rangle$ and another with random initial mixed states, which are distinguished by the color depth of the curve and the error bars. The blue and red curves denote the performances with the observation $\langle\hat{n}_{p}\rangle$ and $\mathcal{I}(t)$, respectively, with error bars. The measurement rate is $\eta=0.5$, and the training and testing time steps are $T=500$.